The line x=y=z-1 intersects the cone x^2+y^2+z^2+2yz+2zx+2xy=0 at exactly one point.
True or false with correct explanation
1
Expert's answer
2020-10-29T18:32:03-0400

"x^2+y^2+z^2+2yz+2zx+2xy=0"


Let x = y = z – 1 = t

"t^2+t^2+(t+1)^2+2t(t+1)+2(t+1)t+2t^2=0"

"4t^2+t^2+2t+1+2t^2+2t+2t^2+2t=0"

"9t^2+6t+1=0"

"(3t+1)^2=0"

"3t+1=0 \implies t=-\frac{1}{3}"


Hence "x=-\frac{1}{3}, y=-\frac{1}{3}, z=\frac{2}{3}"

Therefore "(-\frac{1}{3},-\frac{1}{3},\frac{2}{3})" is the only intersection point, q.e.d. Answer: True.